# Evidence-Focused Fact Summarization for Knowledge-Augmented Zero-Shot Question Answering

### v2

@@ Abstract
Recent studies have investigated utilizing Knowledge Graphs (KGs) to enhance Quesetion Answering (QA) performance of Large Language Models (LLMs), yet structured KG verbalization remains challenging. Existing methods, such as triple-form or free-form textual conversion of triple-form facts, encounter several issues. These include reduced evidence density due to duplicated entities or relationships, and reduced evidence clarity due to an inability to emphasize crucial evidence. To address these issues, we propose EFSum, an Evidence-focused Fact Summarization framework for enhanced QA with knowledge-augmented LLMs. We optimize an open-source LLM as a fact summarizer through distillation and preference alignment. Our extensive experiments show that EFSum improves LLM’s zero-shot QA performance, and it is possible to ensure both the helpfulness and faithfulness of the summary.
@@ 1 Introduction
As one form of external knowledge, Knowledge Graphs (KGs) have been considered as the knowledge source to augment LLMs for enhanced performance in knowledge graph QA (KGQA) Baek et al. (2023a); Wu et al. (2023); Sen et al. (2023). The key challenge of utilizing KGs, which consist of a set of (head entity, relation, tail entity) triples, is to bridge the modality gap between graphs and text. Efforts on connecting the gap mostly fall into either training additional layers to blend two representations, or verbalizing graphs into texts. While training fusion layers Yasunaga et al. (2022a, b) can make models expressive on two different modalities, it takes expensive computaions, and needs to be trained when KGs are updated. On the other hand, recent studies proposed verbalizing KGs into text-form, without restirctions of trainining LLMs for QA. For instance, one strategy is simply concatenating the facts in their triple-form text Baek et al. (2023a), and another entails converting them into semantically-coherent textual description (i.e., free-form text) through the distillation of LLM’s ability to generate text from KGs Wu et al. (2023).
@@
Extensive experiments on two QA benchmark datasets validate the effectiveness of our evidence-focused fact summarization in improving LLM’s zero-shot QA performance. Our approach significantly improves the zero-shot QA performance of LLMs, while also enhancing the helpfulness and faithfulness of the generated fact summaries.
For reproducibility, our codes are publicly available at https://github.com/anon809/EFSum
@@ 2.1 KG-Augmented LLM Prompting for QA
Fact injection for question answering. The last stage is prompting the LLM to generate the answer with the verbalized facts. This process, also known as knowledge-augmented LLM prompting for zero-shot QA, gathers the output as the predicted answer. To handle with insufficient evidence of knowledge, the detailed instructions are provided to allow the LLM to utilize its internal knowledge if needed. The prompts are in Appendix C.2.
@@ 2.2 Analysis on Verbalized Facts
In Figure 2 Upper, it is evident that the number of duplicated tokens is significantly higher in the linearly verbalized texts (i.e., KAPING and Rewrite) compared to the others. This strongly indicates that their outputs contain redundant information stemming from the pre-defined relations within KGs. When considering the ratio of token lengths before and after verbalization (i.e., compression rate), KAPING and Rewrite either maintain or even increase the length, despite maintaining the same amount of information. Consequently, their evidence density remains low or may even decrease during verbalization.
@@ 3.1 LLM Prompting for EFSum
The most straightforward implementation of evidence-focused fact summarization is prompting LLMs to generate a summary given a question and its relevant facts , focusing on the evidence to answer the question. Thus, we present EFSum to verbalize the facts with the help of LLM’s zero-shot capability on summarization; that is, , where is the prompt for summarization. Specifically, we instruct LLMs to turn the input facts into the summary for the scenarios where the summary serves as a context to facilitate QA task. We utilized GPT-3.5-turbo for generating summary of EFSum. The detailed prompts are in Appendix C.2.
@@ 3.2.1 Distillation of Fact Summarization
Reference summary generation. We utilize a closed-source LLM (i.e., GPT-3.5-turbo) to obtain reference fact summaries used for training our summarizer. For each tuple of a question and the relevant facts , we prompt the LLM to transform the set of facts into a concise textual description that highlights the evidence for the question , which is same with EFSum. In the end, we construct the training dataset . The prompt used for summary generation is in Table 5 of Appendix C.2.
@@ 3.2.2 Alignment with Summary Preference
Faithfulness filter. The second filter focuses on the faithfulness of each summary candidate. We use the G-Eval approach Liu et al. (2023) to leverage LLM’s ability to evaluate the consistency between the input facts and the given summary in terms of hallucination. Precisely, the LLM is prompted to examine whether or not the summary contains unfaithful information, which cannot be inferred from the given symbolic facts. Please refer to the detailed prompt in Appendix C.2.
Broad-to-specific paraphrasing. In addition, we adopt the paraphrase process guided by the gold answer . This paraphrasing aims to obtain high-quality summary by refining a broad focus into a specific focus, using the given answer as the main evidence. We prompt the LLM to paraphrase the summary candidate into ; , where is the prompt for paraphrasing. This candidate undergoes another pass through the helpfulness filter and the faithfulness filter. The resulting summary obtained in this manner is much more focused on the QA task than the initially obtained reference summary, called as the broad summary. The detailed prompt is provided in Appendix C.2
@@ 4.2 Effectiveness of Dense Evidence (RQ1)
Compatibility with various retrievers. We investigate the QA performance using different types of fact retrieval, to assess the robustness across various knowledge qualities: randomly selected knowledge (Random), the knowledge possessing the most frequently occurring relation (Popular), and question-relevant knowledge (MPNet). In Table 2, EFSum achieves the highest accuracy across most datasets and QA models, regardless of the retriever used, which implies that our method can exact useful facts from noisy input triples.
@@ 5.2 Knowledge Graph Question Answering
Studies exploring the integration of additional knowledge extracted from KGs, represented as subgraphs, are divided into two main approaches. The first approach utilizes semantic parsing techniques Bao et al. (2016); Luo et al. (2018), which enable the extraction of executable queries from the KG by using contextual information as a parsing reference. Alternatively, the information retrieval approach involves encoding assimilated information using techniques like Graph Neural Networks (GNNs). Several recent studies Yasunaga et al. (2022a, b); Zhang et al. (2022) propose a learning framework that combines GNNs and LMs, allowing concurrent utilization of textual data and KG. In contrast to the conventional KGQA approaches that aim to directly identify the answer entity within KGs, in this work, we mainly focus on how the KG-retrieved facts can be effectively utilized within the zero-shot QA capability of LLMs.
@@ B.2 Various Retrievers
Appendix C Qualitative Examples
C.1 Case Study
@@
In Table 4, we present an example of verbalized facts generated by each method. The example is from Mintaka. EFSum emphasizes evidence necessary to answer the question (i.e., Emilio Estevez’s father is Martin Sheen.) at the beginning, while excluding irrelevant details for brevity, as opposed to other baselines. Duplicated relations (i.e., relations of siblings) are aggregated, and it is clear that our method is more compact than KAPING and Rewrite. Comparing to KG2Text, it also sums up the given triples, but reduces a loss of essential information.
C.2 Prompts
@@
## Caution2: The *summary* should only conatin information of the given triples. |

### v3

@@ Abstract
Recent studies have investigated utilizing Knowledge Graphs (KGs) to enhance Question Answering (QA) performance of Large Language Models (LLMs), yet structured KG verbalization remains challenging. Existing methods, such as triple-form or free-form textual conversion of triple-form facts, encounter several issues. These include reduced evidence density due to duplicated entities or relationships, and reduced evidence clarity due to an inability to emphasize crucial evidence. To address these issues, we propose EFSum, an Evidence-focused Fact Summarization framework for enhanced QA with knowledge-augmented LLMs. We optimize an open-source LLM as a fact summarizer through distillation and preference alignment. Our extensive experiments show that EFSum improves LLM’s zero-shot QA performance, and it is possible to ensure both the helpfulness and faithfulness of the summary.
@@ 1 Introduction
As one form of external knowledge, Knowledge Graphs (KGs) have been considered as the knowledge source to augment LLMs for enhanced performance in knowledge graph QA (KGQA) Baek et al. (2023a); Wu et al. (2023); Sen et al. (2023). The key challenge of utilizing KGs, which consist of a set of (head entity, relation, tail entity) triples, is to bridge the modality gap between graphs and text. Efforts on connecting the gap mostly fall into either training additional layers to blend two representations, or verbalizing graphs into texts. While training fusion layers Yasunaga et al. (2022a, b) can make models expressive on two different modalities, it takes expensive computaions, and needs to be trained when KGs are updated. On the other hand, recent studies proposed verbalizing KGs into text-form, without training LLMs for QA. For instance, one strategy is simply concatenating the facts in their triple-form text Baek et al. (2023a), and another entails converting them into semantically-coherent textual description (i.e., free-form text) through the distillation of LLM’s ability to generate text from KGs Wu et al. (2023).
@@
Extensive experiments on two QA benchmark datasets validate the effectiveness of our evidence-focused fact summarization in improving LLM’s zero-shot QA performance. EFSum outperforms other fact verbalization methods in two key settings: (1) when fixing the token length (density, Section 4.2), and (2) when fixing the number of triples (clarity, Section 4.3) within contextual knowledge. Additionally, our approach enhances the helpfulness and faithfulness of the generated fact summaries. For reproducibility, our codes are publicly available at https://github.com/kk13332488/EFSum.
@@ 2.1 KG-Augmented LLM Prompting for QA
Fact injection for question answering. The last stage is prompting the LLM to generate the answer with the verbalized facts. This process, also known as knowledge-augmented LLM prompting for zero-shot QA, gathers the output as the predicted answer. To handle with insufficient evidence of knowledge, the detailed instructions are provided to allow the LLM to utilize its internal knowledge if needed. The prompts are in Appendix D.2.
@@ 2.2 Analysis on Verbalized Facts
In Figure 2 Upper, it is evident that the number of duplicated tokens is significantly higher in the linearly verbalized texts (i.e., KAPING Baek et al. (2023a) and Rewrite Wu et al. (2023)) compared to the others. This strongly indicates that their outputs contain redundant information stemming from the pre-defined relations within KGs. When considering the ratio of token lengths before and after verbalization (i.e., compression rate), KAPING and Rewrite either maintain or even increase the length, despite maintaining the same amount of information. Consequently, their evidence density remains low or may even decrease during verbalization.
@@ 3.1 LLM Prompting for EFSum
The most straightforward implementation of evidence-focused fact summarization is prompting LLMs to generate a summary given a question and its relevant facts , focusing on the evidence to answer the question. Thus, we present EFSum to verbalize the facts with the help of LLM’s zero-shot capability on summarization; that is, , where is the prompt for summarization. Specifically, we instruct LLMs to turn the input facts into the summary for the scenarios where the summary serves as a context to facilitate QA task. We utilized GPT-3.5-turbo for generating summary of EFSum. The detailed prompts are in Appendix D.2.
@@ 3.2.1 Distillation of Fact Summarization
Reference summary generation. We utilize a closed-source LLM (i.e., GPT-3.5-turbo) to obtain reference fact summaries used for training our summarizer. For each tuple of a question and the relevant facts , we prompt the LLM to transform the set of facts into a concise textual description that highlights the evidence for the question , which is same with EFSum. In the end, we construct the training dataset . The prompt used for summary generation is in Table 5 of Appendix D.2.
@@ 3.2.2 Alignment with Summary Preference
Faithfulness filter. The second filter focuses on the faithfulness of each summary candidate. We use the G-Eval approach Liu et al. (2023) to leverage LLM’s ability to evaluate the consistency between the input facts and the given summary in terms of hallucination. Precisely, the LLM is prompted to examine whether or not the summary contains unfaithful information, which cannot be inferred from the given symbolic facts. Please refer to the detailed prompt in Appendix D.2.
Broad-to-specific paraphrasing. In addition, we adopt the paraphrase process guided by the gold answer . This paraphrasing aims to obtain high-quality summary by refining a broad focus into a specific focus, using the given answer as the main evidence. We prompt the LLM to paraphrase the summary candidate into ; , where is the prompt for paraphrasing. This candidate undergoes another pass through the helpfulness filter and the faithfulness filter. The resulting summary obtained in this manner is much more focused on the QA task than the initially obtained reference summary, called as the broad summary. The detailed prompt is provided in Appendix D.2
@@ 4.2 Effectiveness of Dense Evidence (RQ1)
Compatibility with various retrievers. We investigate the QA performance using different types of fact retrieval, to assess the robustness across various knowledge qualities: randomly selected knowledge (Random), the knowledge possessing the most frequently occurring relation (Popular), and question-relevant knowledge (MPNet). In Table 2, EFSum achieves the highest accuracy across most datasets and QA models, regardless of the retriever used, which implies that our method can extract useful facts from noisy input triples.
@@ 5.2 Knowledge Graph Question Answering
Studies exploring the integration of additional knowledge extracted from KGs, represented as subgraphs, are divided into two main approaches. The first approach utilizes semantic parsing techniques Bao et al. (2016); Luo et al. (2018), which enable the extraction of executable queries from the KG by using contextual information as a parsing reference. Alternatively, the information retrieval approach involves encoding assimilated information using techniques like Graph Neural Networks (GNNs). Several recent studies Yasunaga et al. (2022a, b); Zhang et al. (2022) propose a learning framework that combines GNNs and LMs, allowing concurrent utilization of textual data and KG. Besides, approaches that project text embeddings to graph embeddings Razzhigaev et al. (2023) struggle to avoid hallucinations by LLMs. In contrast to the conventional KGQA approaches that aim to directly identify the answer entity within KGs, in this work, we mainly focus on how the KG-retrieved facts can be effectively utilized within the zero-shot QA capability of LLMs.
@@ 8 Ethical Consideration
9 Acknowledgements
This work was supported by the IITP grants funded by the Korea government (MSIT) (No. RS-2020-II201361; RS-2024-00457882, AI Research Hub Project), and the NRF grant funded by the Korea government (MSIT) (No. RS-2023-00244689).
@@ B.1 Datasets
WebQuestionsSP (WebQSP) Berant et al. (2013); Yih et al. (2016) Semantic Parses Dataset, abbreviated to WebQSP, is a KGQA dataset providing SPARQL queries which allows a direct retrieval from Freebase KG. Due to the cessation of updates to Freebase we adopt Wikidata as the foundational KG. As we mentioned in section 4.1, we utilized WebQSP-WD Sorokin and Gurevych (2018), a dataset that offers questions from WebQSP pre-linked to the Wikidata KG.
@@ B.2 Various Retrievers
Appendix C Additional Experiment Results
C.1 Ablation Study
Dataset | Method | Acc |
WebQSP | EFSum | 0.500 |
w/o Paraphrase | 0.477 | |
w/o Helpfulness | 0.453 | |
w/o All Filters, Parapharase | 0.469 | |
Mintaka | EFSum | 0.338 |
w/o Paraphrase | 0.296 | |
w/o Helpfulness | 0.289 | |
w/o All Filters, Parapharase | 0.299 |
To investigate the effect of different filters and paraphrasing approach on EFSum, in Table 4, we evaluate the performance of EFSum on ablation study. For the experimental setting, we used Flan-T5-XL as a QA model on WebQSP and Mintaka with a maximum token length of contextual knowledge as . Each row indicates whether each filter or paraphrasing approach is used or not. ‘w/o All Filters, Paraphrase’means the model is trained without using any filtering or paraphrasing technique. Table 4 demonstrates that the performance of the distilled model drops filters are removed or the paraphrasing stage is omitted. This result indicates that EFSum shows the best performance when all processes are put together. When the helpfulness filter is removed, some summaries that do not aid the QA task get mixed into the chosen set during DPO training. This disrupts the model’s optimization process, leading to a performance decline. Consequently, removing the helpfulness filter results in the most critical performance drop. The paraphrasing stage also plays a crucial role. Without paraphrasing, the ability to highlight important information within paragraphs decreases. As a result, key information in the summary becomes distracted, making it difficult for the QA model to digest the knowledge. This inevitably leads to a performance decline.
C.2 Generalization Ability
Method | Trained on | Generate Summaries on | Acc |
EFSum | Mintaka | WebQSP | 0.455 |
EFSum | WebQSP | Mintaka | 0.281 |
KAPING | WebQSP | WebQSP | 0.439 |
Rewrite | Mintaka | Mintaka | 0.288 |
We examine the generalization capability of EFSum for unseen datasets. In Table 5, We investigate the cross-dataset experiments on Flan-T5-XL with a maximum token length of contextual knowledge as (i.e. compatible to Table 1). We use EFSum that generates the summary for Mintaka trained on WebQSP, also vice versa. Referring to Table 1, the most superior baselines scores 0.439 and 0.288 for WebQSP and Mintaka, respectively. And EFSum that trained on unseen dataset scores 0.455 and 0.281 for WebQSP and Mintaka, respectively. This result shows that our methods outperform the most of baselines and are competitive to the most effective baselines. This indicates that the proposed evidence-focused approach can effectively summarize the evidence even on unseen datasets.
Appendix D Qualitative Examples
D.1 Case Study
@@
In Table 6, we present an example of verbalized facts generated by each method. EFSum emphasizes evidence necessary to answer the question (i.e., Emilio Estevez’s father is Martin Sheen.) at the beginning, while excluding irrelevant details for brevity, as opposed to other baselines. Duplicated relations (i.e., relations of siblings) are aggregated, and it is clear that our method is more compact than KAPING and Rewrite. Comparing to KG2Text, it also sums up the given triples, but reduces a loss of essential information.
D.2 Prompts
@@
## Caution2: The *summary* should only contain information of the given triples. |